The following units of an item were available for sale during the year

Jan 1. Inventory 30 units @ $130
Mar 10. Purchase 70 units @ $140
Aug. 15. Purchased 10 units @ $145
Dec 10. Purchased 70 units @$150

There are 80 units of the item remaining in the physical inventory on December 31st. The periodic inventory system was used. Determine the inventory cost and cost of merchandise sold using the three main methods
Inventory Method Merchandise Inventory Cost
FIFO (first in, first out)   
LIFO (last in, first out)   
Weighted Avg Cost
Answer #1
Units Unit Cost Total
Jan 1 30 130 3900
Mar 10 70 140 9800
Aug 15 10 145 1450
Dec 10 70 150 10500
Total 180 25650
Average cost 142.50 =25650/180
FIFO (first in, first out):
Merchandise Inventory Cost 11950 =(70*150)+(80-70)*145
Cost of merchandise sold 13700 =25650-11950

LIFO (last in, first out) :

Merchandise Inventory Cost 10900 =(30*130)+(80-30)*140
Cost of merchandise sold 14750 =25650-10900
Weighted Avg Cost:
Merchandise Inventory Cost 11400 =80*142.50
Cost of merchandise sold 14250 =25650-11400